OpenStudy (anonymous):

David measured the distance from his house to a soccer field as 3 km. The actual distance from his house to the soccer field is 2.84 km. What is the approximate percent error in David’s measurement? Round to the nearest tenth of a percent, if necessary. A. 0.16% B. 5.3% C. 5.6% D. 16%

OpenStudy (anonymous):

Let's solve your equation step-by-step. 3− 2.84 2.84 = x 100 Step 1: Simplify both sides of the equation. 2= 1 100 x Step 2: Flip the equation. 1 100 x=2 Step 3: Divide both sides by 1/100. 1 100 x 1 100 = 2 1 100 x=200

OpenStudy (anonymous):

Your answer would be 5.6%

OpenStudy (texaschic101):

(3 - 2.84) / 2.84 = 0.16/2.84 = 0.056 = 5.6%